Given that x-1,2x+5,8x-1 are three consecutive terms of a GP. Find possible value of x and the common ratio of the progression
Answer by richwmiller(17219)

arithmetic
x-1,2x+5,8x-1
(2x+5)-(x-1)=(8x-1)-(2x+5)
x = 12/5
x-1 = 7/5, 2x+5 = 49/5, 8x-1 = 91/5
d=42/5
geometric
(2x+5)/(x-1)=(8x-1)/(2x+5)
(2x+5)/(x-1)=a,b=(8x-1)/(2x+5)
a = -2, b = -2, x = -3/4
ratio is -2 when x=-3/4
a = 3, b = 3, x = 8
ratio is 3 when x=8